A swimmer jumps into a river and swims

straight for the other side at 1.5 km/h [N]

relative to the water. There is a current in the

river of 2.0 km/h [W].

a. What is the swimmer’s velocity relative

to the ground? [2.5 km/h [N53°W]]

b. If the river is 150 m wide, how long will

it take them to cross the river?


1
Expert's answer
2022-01-26T15:30:54-0500

Solution.

"v_r=2.0 km\/h;"

"v_s=1.5 km\/h;"

"a. v=\sqrt{v_r^2+v_s^2};"

"v=\sqrt{4+2.25}=2.5km/h;"

"tan\alpha=\dfrac{2}{1.5}=1.33; \alpha=53^o"


"b.
 L=\dfrac{d}{ cos \alpha};" "d=150m;"


"L=\dfrac{150}{0.6018}=250m;"


"t=\dfrac{L}{v};"


"t=\dfrac{250}{2500}=0.1h;"

Answer:"a. v=2.5km/h; \alpha=53^o;"

"b. t=0.1h."